The sum of three consecutive integers is –27. What are the numbers?
Answer by faceoff57(108)

The sum of three consecutive integers is –27. What are the numbers?
X + (x+1) + (x+2) = -27
X + x + 1 + x + 2 = -27
3x + 3 = -27
3x = -27 - 3
3x = -30
X = -10
-10, -9, -8